Robil can do 2 jobs in 3 days, and Buray can do 5 jobs in 6 days. Robil works for 3 days and then Buray begins to help. How many days will it take to complete a total of 74 jobs?
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

Robil can do 2 jobs in 3 days, and Buray can do 5 jobs in 6 days. Robil works for 3 days and then Buray begins to help. How many days will it take to complete a total of 74 jobs?
Let x = time (days) to perform 74 jobs
then
x(2/3 + 5/6) = 74
multiply both sides by 6:
x(4 + 5) = 444
x(9) = 444
x = 444/9
x = 49.33 days
or
x = 49 days and 8 hours

Answer by stanbon(75887)

Robil can do 2 jobs in 3 days, and Buray can do 5 jobs in 6 days. Robil works for 3 days and then Buray begins to help. How many days will it take to complete a total of 74 jobs?
----
Robil rate = (2/3) job/day
Buray rate = (5/6) job/day
--------------------------------
Together rate: 1/x job/day
Equation:
rate + rate = together rate
2/3 + 5/6 = 1/x
(12+15)/18 = 1/x
27/18 = 1/x
3/2 = 1/x
Together rate: 3/2 job/day
---------------------------------
Robil works for 3 days and then Buray begins to help. How many days will it take to complete a total of 74 jobs?
----
In 3 days Buray finishes 3(5/6) = (5/2) jobs
That leaves 71 1/2 jobs to do together
-----------------
Equation:
x(3/2) = 71 1/2
(3/2)x = 143/2
x = (2/3)(143/2)
x = 143/3 = 45 3/4 days
================================
Cheers,
Stan H.
============